### v1

@@ 3.1 Pre-Training Hessian-Free Laplace
for observation variance hyperparameter $\sigma^{2}$. Under the common Gaussian likelihood assumption, Eq. 29 implies that pre-trained HFL is performed by data augmentation, without requiring any adjustment to the model architecture or training procedure.
Parameter Uncertainty Quantification
The same pre-training approach also enables parameter uncertainty quantification by finding the regularizer $\hat{t}$ that minimizes the average squared L2 norm over the $K$ dimensions of parameters,
Theorem 3.2 applies to Eq. 30 in combination with absolute value regularization to provide the pre-training objective,
which corresponds to a set of sparsity-inducing Laplace priors over the weights. Thus, when used in concert with a Laplace log prior $\log p(\theta)$, the regularization term for HFL is absorbed, with $\lambda$ perturbing the precision parameter. In this case, it is particularly convenient to implement HFL for parameter uncertainty, as one needs only to adjust the strength of the existing regularizer and use,
$\hat{\sigma}_{\theta}^{2}(x):=\frac{1}{\lambda}|\hat{\theta}^{(\hat{r},\lambda)}-\hat{\theta}|,$ (32)
as the parameter variance.
4 Related Work
The Laplace approximation was first formulated in BDL for small-scale neural networks MacKay (1992). Since then, larger architectures have necessitated Hessian approximations, including Gauss-Newton Foresee & Hagan (1997) and Kronecker factorization that places a block diagonal structure on the covariate in line with assumed independence between the layers of a network Martens & Grosse (2015).
In recent years, with the aforementioned scalable approximations and renewed appreciation of the advantages of Laplace as a post-hoc uncertainty method Daxberger et al. (2021), a number of new directions were addressed, including relating neural network loss functions to Gaussian process inference Khan et al. (2019) and exploring the advantages of locally linearized Laplace in out-of-distribution evaluation Foong et al. (2019). Underfitting in Laplace Lawrence (2001) is mitigated in the GGN approximation by switching to the generalized linear model implied by the local linearization Immer et al. (2021).
Owing to the ubiquity of the Hessian of the objective and its computational demands, approaches to avoiding its direct evaluation are developed in areas of research distinct of Laplace. In particular, Hessian-free optimization uses the conjugate gradient method to iteratively update a matrix-vector product as part of an inner loop in Newton’s method second-order optimization Martens & Sutskever (2011); Pearlmutter (1994).
@@ 4 Related Work
HFL (this paper) PICP: 0.9375 PICP: 0.9333 PICP: 0.8750 PICP: 0.8444 PICP: 1.0000 PICP: 0.8375 PICP: 1.0000 PICP: 0.9250
@@
Table 1: Uncertainty quantification performance across datasets and methods. Top performing metrics, as defined in Section 5, are in bold, with ties broken by the other metrics. Full Hessian is a benchmark comparison that is intractable at scale.
(a) Full Hessian
(b) GGN
(c) Eigen. Approx.
(d) HFL (this paper)
5 Experiments
Figure 3: The uncertainty ranges of a random subset of 30 parameters of the neural network indicate that many parameters include 0 in their intervals. HFL recovers much of the structure of these uncertainties via additional regularization of the network parameters in the objective function.
@@ 5 Experiments
We explore four data generating distributions designed to test both in-distribution inference and extrapolation to “in-between” uncertainty (Foong et al., 2019),
A feedforward neural network is used with 50 hidden units and $\tanh$ activations to ensure that the Hessian exists, and optimized using Adam (Kingma & Ba, 2017) with a learning rate of 0.005. Further experimental details are given in Appendix A. Four methods to quantifying the epistemic uncertainty of the network are compared,
  • GGN using Eq. 8 with only the second term in likelihood Hessian (Eq. 9).
  • HFL (this paper), using Algorithm 2.
The metrics used to evaluate uncertainty performance are,
The results of the evaluations are summarized in Table 1. We find that HFL performs commensurately with other approximations, and indeed the full Hessian, despite taking a significantly different approach. In particular, in several cases it achieves the best out-of-distribution (OOD) performance, likely due to its ability to avoid the GGN assumption that ignores an important part of the curvature around the MAP. To illustrate this point, we plot the predicted mean output for the methods on the Sin-Inbetween OOD task in Figure. 2. There is a noticeable difference to the GGN and eigenvector approximation, implying that in practice the residual of the loss may not be close to zero.
We also evaluated the parameter regularization approach given in Eq. 31. Figure 3 shows the 1 standard deviation interval about the predicted mean under the various methods. Recall that HFL for parameter uncertainty requires only to apply a small amount of additional L1 regularization to the parameters. Nonetheless, the intervals indicated often align with those requiring full Hessian inversion. These parameter uncertainties may be used for model introspection. As an example, we considered the task of adding sparsity to a network, setting parameters to 0 when their uncertainty intervals overlap with 0. The mean absolute errors of the resulting predictions are shown in Figure 4. HFL does competitively, particularly for in-between uncertainty. An outlier result is the absolute error of the full Hessian in the Quadratic-Inbetween in-distribution task. This was due to mis-selecting an entire group of units to be set to 0, greatly impacting the shape of the predicted function.
Limitations
HFL is subject to the same assumptions and limitations as the Laplace approximation, primarily, that it is a local method centered around the optimum of the joint distribution and the curvature is evaluated at one location (ignoring any other modes). In practice, early stopping and highly overparameterized networks break these assumptions, reducing the quality of the uncertainties. In addition, pre-trained HFL introduces error in the set of evaluation points embedded in the regularized objective, though this has so far been observed to be mild in practice.
6 Conclusions & Future Work
In this paper we developed HFL, a set of approaches to Laplace approximation without explicit Hessian evaluation and inversion. Instead, two point estimates are used, the usual MAP estimate and a regularized form of the MAP. Theoretical and experimental evidence supports this method both in terms of accuracy of predictive variance and the favorable scaling properties in the number of parameters of the model. Future work is required to investigate the role of these approximations in architectures of greater complexity as well as their potential role in directing exploration, experimentation, training data selection, and model introspection.
Appendix A Further Experimental Details
Figure 5: Hyperparameter sweeps for observational variance.
For both the quadratic and sin datasets, the observational variance hyperparameter was selected through grid search over the set $\{0.005,0.01,0.05,0.1,0.5,1.0\}$ minimizing negative log likelihood for a held-out validation set of equal size to the training dataset. The results are summarized in Figure 5, with the optimal hyperparameter chosen indicated by a marker. The prior variance was fixed at 3.0, which was the minimum value to the nearest 0.1 for which the Hessian was sufficiently conditioned across all tasks to allow for inversion.

### v2

@@ B.2 Data-Augmentation Objective
for observation variance hyperparameter $\sigma^{2}$. Under the common Gaussian likelihood assumption, Eq. 28 implies that pre-trained RegVar is performed by data augmentation, without requiring any adjustment to the model architecture or training procedure.
Algorithm 3 Amortized RegVar for Parameter Uncertainty
  Input: network $f_{\theta}$, fitted MAP $\hat{\theta}$, training data $\mathcal{D}$ of size $n$, evaluation inputs $\{\hat{x}_{i=1}^{m}\}$, scalar $\lambda$, step size schedule $\gamma(\cdot)$
  Warm start $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}$
  Initialize step counter $j=0$
  repeat
     Sample training example $(x_{i},y_{i})\sim\mathcal{D}$
     Follow gradient $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}^{(\hat{r},\lambda)}+n\gamma(j)\nabla_{\theta}\mathcal{L}_{i}^{\mathrm{(PU)}}\;\;\;$ //See Eq. 18
     $j\leftarrow j+1$
  until $\hat{\theta}^{(\hat{r},\lambda)}$ converges
  Return: parameter variance $\hat{\sigma}_{\theta}^{2}(x):=\frac{1}{\lambda}|\hat{\theta}^{(\hat{r},\lambda)}-\hat{\theta}|$
B.3 RegVar for Parameter Uncertainty
For completeness, we present Algorithm 3 that corresponds to the parameter uncertainty objective given in Eq. 18 in the main text. We apply this algorithm in Section C.2.
Appendix C FURTHER EXPERIMENTS AND HYPERPARAMETER SELECTION
In this section, we evaluate RegVar and compare against the full Hessian and its approximations.
C.1 Synthetic Data
We explore four data generating distributions designed to test both in-distribution inference and extrapolation to “in-between” uncertainty (Foong et al., 2019). A feedforward neural network is used with 50 hidden units and $\tanh$ activations to ensure that the Hessian exists, and optimized using Adam (Kingma and Ba, 2017) with a learning rate of 0.005. Four methods for quantifying the epistemic uncertainty of the network are compared,
  • GGN using Eq. 4 with only the second term in likelihood Hessian (Eq. 5).
  • RegVar (this paper), using Algorithm 1.
The metrics used to evaluate uncertainty performance are,
(a) Full Hessian
(b) GGN
(c) Eigen. Approx.
(d) HFL (this paper)
The synthetic data-generating distributions are,
(a) Hyperparameter sweeps for observational variance.
(b) Lambda sweep
Figure 7: Hyperparameter selection for modeling and RegVar algorithm.
@@ C.1 Synthetic Data
RegVar PICP: 0.9375 PICP: 0.9333 PICP: 0.8750 PICP: 0.8444 PICP: 1.0000 PICP: 0.8375 PICP: 1.0000 PICP: 0.9250
@@
Table 3: Uncertainty quantification performance across datasets and methods. Full Hessian (top line) is intractable at scale and included here for reference.
For both the quadratic and sin datasets, the observational variance hyperparameter was selected through grid search over the set $\{0.005,0.01,0.05,0.1,0.5,1.0\}$ minimizing negative log likelihood for a held-out validation set of equal size to the training dataset. The results are summarized in Figure 7(a), with the optimal hyperparameter chosen indicated by a marker. The prior variance was fixed at 3.0, which was the minimum value to the nearest 0.1 for which the Hessian was sufficiently conditioned across all tasks to allow for inversion.
Figure 8: The uncertainty ranges of a random subset of 30 parameters of the neural network indicate that many parameters include 0 in their intervals. RegVar recovers much of their structure just via additional regularization of the network parameters in the objective function.
@@
The results of the evaluations are summarized in Table 3. We find that RegVar performs commensurately with other approximations, and indeed the full Hessian, despite taking a significantly different approach. In particular, in several cases it achieves the best out-of-distribution (OOD) performance, likely due to its ability to avoid the GGN assumption that ignores an important part of the curvature around the MAP. To illustrate this point, we plot the predicted mean output for the methods on the Sin-Inbetween OOD task in Figure. 6. There is a noticeable difference to the GGN and eigenvector approximation, implying that in practice the residual of the loss may not be close to zero.
C.2 Parameter Uncertainty Experiment
We also evaluated the parameter regularization approach given in Eq. 18. Figure 8 shows the 1 standard deviation interval about the predicted mean under the various methods. Recall that RegVar for parameter uncertainty requires only to apply a small amount of additional L1 regularization to the parameters. Nonetheless, the intervals indicated often align with those requiring full Hessian inversion. These parameter uncertainties may be used for model introspection. As an example, we considered the task of adding sparsity to a network, setting parameters to 0 when their uncertainty intervals overlap with 0. The mean absolute errors of the resulting predictions are shown in Figure 9. RegVar does competitively, particularly for in-between uncertainty. An outlier result is the absolute error of the full Hessian in the Quadratic-Inbetween in-distribution task. This was due to mis-selecting an entire group of units to be set to 0, greatly impacting the shape of predictions.